# Holographic Generalised Gukov-Witten Defects

### v1

@@ 1   Introduction
Spindle and disc solutions have received significant attention in recent literature. See [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] for a necessarily incomplete set of references. These solutions present a generalisation of the usual lore of lower dimensional supergravity by allowing for certain (mild) singularities. In particular, they include conical defects. Unlike in [25] where similar conical defects were introduced on higher genus Riemann surfaces, the spindle and disc represent respectively a sphere with (one or) two conical defects at the poles, or a topological disc with one defect at the centre. See Figure 1 for a graphical representation. Provided that the orders of the conical defects are co-prime, these are so-called bad orbifold [26], which do not admit a constant curvature metric. For this reason, such solutions preserve supersymmetry in a novel way. In fact for the spindle there are often two options, dubbed the twist or anti-twist [15].
@@
The main claim of this work, in line with the claims above, is that such an interpretation is valid more generally. In particular, we will argue that the conical defects in the five-dimensional spindle and disc solutions correspond to a class of generalised Gukov-Witten defects [28, 29]. To argue for these claims we will zoom in to a single conical defect of the spindle and study a class of holographic defect solutions. Similar defect solutions have been considered in gauged supergravity in [30, 31, 32, 33, 34, 35] while in ten- or eleven-dimensional supergravity several classes of $\frac{1}{2}$-BPS defects have been discussed as bubbling defect solutions [36, 37, 38, 39, 40, 41, 42, 43]. In this work we generalise the analysis of said papers by starting from a class of gauged supergravity solutions with a conical defect in the bulk. Unlike [30, 34] this generalisation allows us to holographically access a class of $\frac{1}{2}$-BPS defects and significantly enlarge the class of accessible, less supersymmetric, defects.
@@
$T^{\mu}_{\mu}\bigg{|}_{\Sigma}=\frac{1}{24\pi}\left(bR_{\Sigma}+d_{1}\Pi^{2}+d_{2}W^{ab}{}_{ab}\right)+\cdots$ (1.2)
Where $R_{\Sigma}$ is the Ricci curvature of the defect world-volume, $\Pi^{2}$ is a particular contraction of the extrinsic curvature and $W^{ab}{}_{ab}$ is the contraction of the bulk Weyl tensor, pulled back to the defect world volume. In this formula $b$ is the only type A anomaly. For the $\frac{1}{2}$-BPS defects, the expressions above agree precisely with the expected defect anomalies for (axially symmetric) Gukov-Witten defects [38, 40, 41] while for their less supersymmetric cousins these provide explicit predictions from holography.
@@
This difference is reminiscent of the story proposed in [54], where two alternative prescriptions to describe surface defects were proposed. On the one hand, surface defects are described by specifying the singularities along a co-dimension two locus, similar as the disorder defects discussed in this work, while on the other hand they show that there is an equivalent description of these defects by including conical defects in space-time. It would be interesting to understand the connection of this work with our setup to make the above story more precise. A starting point to understand this setup would be to study the spindle index discussed in [55, 56]. In these works ABJM is localised on $S^{1}\times\Sigma$, giving rise to an index that closely resembles the standard index on $S^{1}\times S^{2}$ with the insertion of vortex lines at the poles of the $S^{2}$.
@@ 2.1   Defect observables
where we always take the operator to lie at the boundary of ${\mathop{\rm AdS}\nolimits}_{3}\times S^{3}$ and normalise the one-point function so that $\left\langle{1}\right\rangle_{V}=1$.
In this work we will be particularly interested in the one-point function of the stress-tensor $T_{\mu\nu}$. As shown in [38] this one-point function takes the form
@@
$\left\langle{T^{\mu}{}_{\mu}}\right\rangle\supset\delta(\Sigma)\frac{1}{24\pi}(b\bar{R}+d_{1}{\Pi}^{2}+d_{2}W_{ab}{}^{ab})\,.$ (2.9)
@@ 4.3   Comparison with bubbling surface operator backgrounds
$\partial_{i}\partial_{i}a+y\partial_{y}\left(\frac{\partial_{y}a}{y}\right)=\sum_{k=1}^{M}2\pi y_{k}\,\delta(y-y_{k})\delta^{2}(\vec{x}-\vec{x_{k}})\,,$ (4.15)
@@
$a_{k}(x_{1},x_{2},y)=\frac{(\vec{x}-\vec{x_{k}})^{2}+y^{2}+y_{k}^{2}}{2\sqrt{((\vec{x}-\vec{x_{k}})^{2}+y^{2}+y_{k}^{2})^{2}-4y_{k}^{2}y^{2}}}-\frac{1}{2}\,.$ (4.17)
As shown in [38, 41] such backgrounds can be interpreted as holographic surface defects where the positions $(y_{k},\vec{x_{k}})$ appearing as sources on the right-hand side are related to the parameters $(\beta_{k},\gamma_{k})$ and to the choice of Levi subgroup $L=\prod_{k=1}^{M}U(N_{k})$. More precisely, we have
$y_{k}^{2}=\frac{N_{k}}{N},\quad\vec{x_{k}}=2\pi(\beta_{k},\gamma_{k})$ (4.18)
Moreover, the parameters $\alpha_{k}$, $\eta_{k}$ can be related to the periods of the 2-form potentials $B_{2}$ and $C_{2}$ of Type IIB supergravity,
$\alpha_{k}=-\frac{1}{2\pi}\int_{\mathcal{C}_{k}}B_{2}\,,\qquad\eta_{k}=-\frac{1}{2\pi}\int_{\mathcal{C}_{k}}C_{2}\,,$ (4.19)
where $\mathcal{C}_{l}$ are non-trivial two-cycles constructed by fibring an $S^{1}$ over the interval connecting $y\in[y_{k},\infty)$.
@@
$M=k\,,\qquad y_{k}^{2}=\frac{N_{k}}{N}=(1-q_{1})=\frac{1}{k}\,,\qquad\vec{x_{k}}=2\pi(\beta_{k},\gamma_{k})=0\,.$ (4.24)
@@
$\alpha_{k}=\alpha_{1}\,.$ (4.25)
This concludes the identifications and therefore allows us to claim a clear interpretation of the defect solutions as holographic Gukov-Witten surface defects. With non-vanishing $Y_{I}$ the axial symmetry is broken, potentially allowing non-zero values of $\beta_{k}$ and $\gamma_{k}$. However, the map to LLM coordinates in this case is much more complicated and we leave a full analysis for future work. A preliminary analysis indicates that five-dimensional supergravity solutions are not fine-grained enough to give rise to fully localised (non-axially symmetric) sources in ten dimensions.
@@ 5.1   Symmetries and anomalies
at the points where the $S^{1}$ fibre shrinks the fibration has a monopole source. The solutions we found correspond to a single monopole source at $(y,x_{i})=(\sqrt{w_{0}},0)$, which we can interpret as coordinates on this three-dimensional space. More generally, the bubbling surface defect solutions of [38, 40], have $p$ monopole sources at positions $(y_{a},x_{i,a})$, $a=1,\dots,p$, where flux quantisation imposes $\sum_{a}Ny_{a}k_{a}=N$, with $Ny_{a}$ and $k_{a}$ integers. At the locations of the monopole source, the 4d space is locally $\mathbf{R}^{4}/\mathbf{Z}_{k_{a}}$. Hence, we immediately find the defect anomaly polynomial corresponding to the $\frac{1}{2}$-BPS defects considered in this work as [52],
$\mathcal{I}_{4}=-\chi\left(\mathrm{SO}(4)\right)\sum_{a=1}^{p}l_{a}\left(N^{2}y_{a}^{2}-Ny_{a-1}\right)\,,\qquad l_{a}=\sum_{b=a}^{p}k_{b}\,.$ (5.8)
@@
$b_{R}=b_{L}=$ $\,-3\sum_{a=1}^{p}l_{a}\left(y_{a}^{2}-y_{a-1}\right)$ (5.10)
$=$ $\,-3N^{2}\left(1-\frac{1}{k}\right)\,.$ (5.11)
@@ Acknowledgements
It is with pleasure that we thank Chris Couzens, Federico Bonetti and Brandon Robinson for useful and inspiring discussions. In particular, we thank Chris Couzens for comments on the draft. The contributions of PB were made possible through the support of grant No. 494786 from the Simons Foundation (Simons Collaboration on the Non-perturbative Bootstrap) and the ERC Consolidator Grant No. 864828, titled “Algebraic Foundations of Supersymmetric Quantum Field Theory” (SCFTAlg). LT is supported by a STFC Doctoral Training Partnership grant.

### v2

@@ 1   Introduction
Spindle and disc solutions have received significant attention in recent literature. See [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] for a necessarily incomplete set of references. These solutions present a generalisation of the usual lore of lower dimensional supergravity by allowing for certain (mild) singularities. In particular, they include conical defects. Unlike in [25] where similar conical defects were introduced on higher genus Riemann surfaces, the spindle and disc represent respectively a sphere with (one or) two conical defects at the poles, or a topological disc with one defect at the centre. See Figure 1 for a graphical representation. Provided that the orders of the conical defects are co-prime, these are so-called bad orbifold [26], which do not admit a constant curvature metric. For this reason, such solutions preserve supersymmetry in a novel way. In fact for the spindle solutions there are often two options, dubbed the twist and anti-twist [15].
@@
The main claim of this work, in line with the claims above, is that such an interpretation is valid more generally. In particular, we will argue that the conical defects in the five-dimensional spindle and disc solutions correspond to a class of generalised Gukov-Witten defects [28, 29]. To argue for these claims we will zoom in to a single conical defect of the spindle and study a class of holographic defect solutions. Similar defect solutions have been considered in gauged supergravity in [30, 31, 32, 33, 34, 35] while in ten- or eleven-dimensional supergravity several classes of $\frac{1}{2}$-BPS defects have been discussed as bubbling defect solutions [36, 37, 38, 39, 40, 41, 42, 43, 44]. In this work we generalise the analysis of said papers by starting from a class of gauged supergravity solutions with a conical defect in the bulk. Unlike [30, 34] this generalisation allows us to holographically access a class of $\frac{1}{2}$-BPS defects and significantly enlarge the class of accessible, less supersymmetric, defects.
@@
$T^{\mu}_{\mu}\bigg{|}_{\Sigma}=-\frac{1}{24\pi}\left(bR_{\Sigma}+d_{1}\Pi^{2}-d_{2}W^{ab}{}_{ab}\right)+\cdots$ (1.2)
Where $R_{\Sigma}$ is the Ricci curvature of the defect world-volume, $\Pi^{2}$ is a particular contraction of the extrinsic curvature and $W^{ab}{}_{ab}$ is the contraction of the bulk Weyl tensor, pulled back to the defect world volume. In this formula $b$ is the only type A anomaly. For the $\frac{1}{2}$-BPS defects, the expressions above agree precisely with the expected defect anomalies for (axially symmetric) Gukov-Witten defects [38, 40, 41, 44] while for their less supersymmetric cousins these provide explicit predictions from holography.
@@
This difference is reminiscent of the story proposed in [55], where two alternative prescriptions to describe surface defects were proposed. On the one hand, surface defects are described by specifying the singularities along a co-dimension two locus, similar as the disorder defects discussed in this work, while on the other hand they show that there is an equivalent description of these defects by including conical defects in space-time. It would be interesting to understand the connection of this work with our setup to make the above story more precise. A starting point to understand this setup would be to study the spindle index discussed in [56, 57, 58]. In these works ABJM is localised on $S^{1}\times\Sigma$, giving rise to an index that closely resembles the standard index on $S^{1}\times S^{2}$ with the insertion of vortex lines at the poles of the $S^{2}$.
@@ 2.1   Defect observables
where we always take the operator to lie at the boundary of ${\mathop{\rm AdS}\nolimits}_{3}\times S^{1}$ and normalise the one-point function so that $\left\langle{1}\right\rangle_{V}=1$.
In this work we will be particularly interested in the one-point function of the stress-tensor $T_{\mu\nu}$. As shown in [38, 44] this one-point function takes the form
@@
$\left\langle{T^{\mu}{}_{\mu}}\right\rangle\supset-\delta(\Sigma)\frac{1}{24\pi}(b\bar{R}+d_{1}{\Pi}^{2}-d_{2}W_{ab}{}^{ab})\,.$ (2.9)
@@ 4.3   Comparison with bubbling surface operator backgrounds
$\partial_{i}\partial_{i}a+y\partial_{y}\left(\frac{\partial_{y}a}{y}\right)=\sum_{l=1}^{M}2\pi y_{l}\,\delta(y-y_{l})\delta^{2}(\vec{x}-\vec{x_{l}})\,,$ (4.15)
@@
$a_{l}(x_{1},x_{2},y)=\frac{(\vec{x}-\vec{x_{l}})^{2}+y^{2}+y_{l}^{2}}{2\sqrt{((\vec{x}-\vec{x_{l}})^{2}+y^{2}+y_{l}^{2})^{2}-4y_{l}^{2}y^{2}}}-\frac{1}{2}\,.$ (4.17)
As shown in [38, 41] such backgrounds can be interpreted as holographic surface defects where the positions $(y_{l},\vec{x_{l}})$ appearing as sources on the right-hand side are related to the parameters $(\beta_{l},\gamma_{l})$ and to the choice of Levi subgroup $L=\prod_{l=1}^{M}U(N_{l})$. More precisely, we have
$y_{l}^{2}=\frac{N_{l}}{N},\quad\vec{x_{l}}=2\pi(\beta_{l},\gamma_{l})$ (4.18)
Moreover, the parameters $\alpha_{l}$, $\eta_{l}$ can be related to the periods of the 2-form potentials $B_{2}$ and $C_{2}$ of Type IIB supergravity,
$\alpha_{l}=-\frac{1}{2\pi}\int_{\mathcal{C}_{l}}B_{2}\,,\qquad\eta_{l}=-\frac{1}{2\pi}\int_{\mathcal{C}_{l}}C_{2}\,,$ (4.19)
where $\mathcal{C}_{l}$ are non-trivial two-cycles constructed by fibring an $S^{1}$ over the interval connecting $y\in[y_{l},\infty)$.
@@
$M=k\,,\qquad y_{l}^{2}=\frac{N_{l}}{N}=(1-q_{1})=\frac{1}{k}\,,\qquad\vec{x_{l}}=2\pi(\beta_{l},\gamma_{l})=0\,.$ (4.24)
@@
$\alpha_{l}=\alpha_{1}\,.$ (4.25)
This concludes the identifications and therefore allows us to claim a clear interpretation of the defect solutions as holographic Gukov-Witten surface defects. With non-vanishing $Y_{I}$ the axial symmetry is broken, potentially allowing non-zero values of $\beta_{l}$ and $\gamma_{l}$. However, the map to LLM coordinates in this case is much more complicated and we leave a full analysis for future work. A preliminary analysis indicates that five-dimensional supergravity solutions are not fine-grained enough to give rise to fully localised (non-axially symmetric) sources in ten dimensions.
@@ 5.1   Symmetries and anomalies
at the points where the $S^{1}$ fibre shrinks the fibration has a monopole source. The solutions we found correspond to $M=k$ coincident monopole sources at $(y,x_{i})=(\sqrt{w_{0}},0)$, which we can interpret as coordinates on this three-dimensional space. More generally, the bubbling surface defect solutions of [38, 40], have $p$ monopole sources at positions $(y_{a},x_{i,a})$, $a=1,\dots,p$, where flux quantisation imposes $\sum_{a}Ny_{a}k_{a}=N$, with $Ny_{a}$ and $k_{a}$ integers. At the locations of the monopole source, the 4d space is locally $\mathbf{R}^{4}/\mathbf{Z}_{k_{a}}$. Hence, we immediately find the defect anomaly polynomial corresponding to the $\frac{1}{2}$-BPS defects considered in this work as [53],
$\mathcal{I}_{4}^{\rm defect}=-\chi\left(\mathrm{SO}(4)\right)\sum_{a=1}^{p}l_{a}\left(N^{2}y_{a}^{2}-Ny_{a-1}\right)\,,\qquad l_{a}=\sum_{b=a}^{p}k_{b}\,.$ (5.8)
@@
$b_{R}=b_{L}=$  $\,3\sum_{a=1}^{p}l_{a}\left(y_{a}^{2}-y_{a-1}\right)$ (5.10)
$=$  $\,3N^{2}\left(1-\frac{1}{k}\right)\,.$ (5.11)
@@ 5.2   Renormalised Action
The partition function in the presence of a defect scales as
$\left\langle{1}\right\rangle_{\Sigma}=R^{b/3}\,,$ (5.19)
where $R$ here is the ${\mathop{\rm AdS}\nolimits}_{3}$ length scale [40, 44]. Comparing this with our on-shell action we have
$\left\langle{1}\right\rangle_{\Sigma}=\mathrm{e}^{-S_{\rm defect}}=R^{\frac{\pi}{2G_{N}}(1-w_{0})}\,,$ (5.20)
where we used the regularised volume $\mathbf{V}_{{\mathop{\rm AdS}\nolimits}_{3}}=-2\pi\log R$. As we will see below this matches the expected value of $b$.
@@ Acknowledgements
It is with pleasure that we thank Chris Couzens, Federico Bonetti, Jaume Gomis and Brandon Robinson for useful and inspiring discussions. In particular, we thank Chris Couzens for comments on the draft. The contributions of PB were made possible through the support of grant No. 494786 from the Simons Foundation (Simons Collaboration on the Non-perturbative Bootstrap) and the ERC Consolidator Grant No. 864828, titled “Algebraic Foundations of Supersymmetric Quantum Field Theory” (SCFTAlg). LT is supported by a STFC Doctoral Training Partnership grant.